OpenStudy (unklerhaukus):

if \(f=f(x,y)\) the total differential is \[\text df= \frac{\partial f}{\partial x}\text dx+\frac{\partial f}{\partial y}\text dy\] right?

OpenStudy (unklerhaukus):

or is the total differential \[\frac{\text df}{\text dx}= \frac{\partial f}{\partial x}+\frac{\partial f}{\partial y}\frac{\text dy}{\text dy}\]

OpenStudy (experimentx):

\[ (xyz) = c \\ \implies yz dx + xz dy + xy dz = 0\]

OpenStudy (unklerhaukus):

whta?

OpenStudy (experimentx):

yes it is as what you said.

OpenStudy (anonymous):

the first

OpenStudy (unklerhaukus):

is the total differential \(\text df\) or \(\frac{\text df}{\text dx}\),

OpenStudy (unklerhaukus):

ah ok, the total differential is just \[\text df\] that's all i was checking

OpenStudy (unklerhaukus):

thanks

OpenStudy (unklerhaukus):

does the other thing have a name ?

OpenStudy (anonymous):

i think so let me check

OpenStudy (unklerhaukus):

the total differential in the x direction must have a better name

OpenStudy (anonymous):

the second one you did looks sorta like teh chain rule for partial derivatives.. but i'm not sure

OpenStudy (unklerhaukus):

ops typo \[\frac{\text df}{\text dx}= \frac{\partial f}{\partial x}+\frac{\partial f}{\partial y}\frac{\text dy}{\text dx}\]

OpenStudy (unklerhaukus):

i suppose \( \frac{\text df}{\text dx}\) is the total derivative ,

OpenStudy (anonymous):

yeah but that's not the actual equation correct?

OpenStudy (unklerhaukus):

the question was \[f(x,y)=y^3(1-x^2)-\cos^2x-3\], find the total differential

OpenStudy (unklerhaukus):

but i think i can do it

OpenStudy (anonymous):

yeah that's for that but the total derivative for an f(x,y) would be \[\frac{dz}{dt}=\frac{\delta f}{\delta x}\frac{dx}{dt}+\frac{\delta f}{\delta y}\frac{dy}{dt}\]

OpenStudy (unklerhaukus):

\partial 's ?

OpenStudy (unklerhaukus):

\[f(x,y)=y^3(1-x^2)-\cos^2x-3\] \[\frac{\partial f}{\partial x}=\frac{\partial }{\partial x}\left(y^3(1-x^2)-\cos^2x-3\right)=-2xy^3+2\sin x\cos x\]\[\frac{\partial f}{\partial y}=\frac{\partial }{\partial y}\left(y^3(1-x^2)-\cos^2x-3\right)=3y^2(1-x^2)\] \[\text df= \frac{\partial f}{\partial x}\text dx+\frac{\partial f}{\partial y}\text dy\]\[\text df=\left(-2xy^3+2\sin x\cos x\right)\text dx+\left(3y^2(1-x^2)\right) \text dy\]

OpenStudy (unklerhaukus):

if the queston had asked fot the total derivavative \[\frac{\text df}{\text dx}=\left(-2xy^3+2\sin x\cos x\right)+\left(3y^2(1-x^2)\right) \frac{\text dy}{\text dx}\]

OpenStudy (experimentx):

getting differentiation equation out of the given equation is not that difficult ... getting the surface out of differential is challenging.

OpenStudy (unklerhaukus):

how'd ya do that exactly?

OpenStudy (experimentx):

there are bunch of ways ... if one doesn't work then other will ...(if satisfies the condition of integrability) ... easiest one is inspection. \[ yzdx+xzdy+xydz=0 \] this will give \[ xyz = c\]

OpenStudy (unklerhaukus):

some kinda hyperbolic something?